elation ategory 2 company capial) Compute the cost for the following sources of financing. (a) Debt that has a $100 par value (face value) and a contractor coupon interest rate of 11% retained prohts if the company tax rates 30%. 14-1(dividual or component costs of ssume a company tax rate of 30%. A new issue would have issue costs of 5% of the $112.50 market value. The debt matures in 10 years (b) A new ordinary share issue that paid an 18 cents dividend last year. Eanings per share have grown at a rate of 7% per year. This growth rate is expected to continue into the foresee- able future The company maintains a constant dividend earnings ratio of 30%, The pnce of this share is now $2.75, but 5% issue costs are anticipated. Dividends are fully franked (for a taxation category I company). (c) Internal equity where the current market price of the ordinary shares is $4.30. The expected dividend this year should be 35 cents, increasing thereafter at a 7% annual growth rate. Dividends are fully franked (for a taxation category I company) 13.5 cents per share. If a new issue is offered,issue costs will be 12% of the current price of $1.75. (d) Preference shares for a taxation category 2 company paying an unfranked dividend of
Answer #1

Part 1

Issue or flotation costs = 5%

Therefore, new price = 112.50*(1-5%) = 106.875

Coupon payment (PMT) = 100*11% = 11

Now apply excel formula “RATE” and its parameters will be

NPER = 10

PMT = 11

PV =-106.875

FV = 100

Now apply formula

=RATE(nper, pmt, pv, [fv])

=RATE(10,11,-106.875,100)

= 9.89%

Cost before tax = 9.89%

Cost after tax = 9.89%*(1-t) = 9.89%*(1-30%) = 6.92%

Part 2

Cost of equity = D1/(P*(1-f) + g

D1 = D0*(1+g)

F = issue cost or flotation cost

G = growth rate

P = market price

= ((0.18*1.07)/((2.75*(1-5%)))+7%

= 14.37%

Part 3

Cost of internal common equity = D1/ P + g

= (0.35/4.3) + 7%

= 15.14%

Part 4

Cost of preference shares = D1 / (p*(1-f)

= 0.135 / (1.75*(1-12%))

= 8.77%